# Spectral flow for minimal $W$-algebras and application to unitarity of their representations

### v1

@@ Abstract.
Using spectral flow, we provide a proof of [9, Theorem 9.17] on unitarity of Ramond twisted non-extremal representations of minimal $W$-algebras that does not rely on the still conjectural exactness of the twisted quantum reduction functor (see Conjecture 9.11 of [9]). When $\mathfrak{g}=spo(2|2n)$, $F(4)$, $D(2,1;\frac{m}{n})$, it is also proven that the unitarity of extremal (=massless) representations of the minimal $W$-algebra $W^{k}_{\min}(\mathfrak{g})$ in the Ramond sector is equivalent to the unitarity of extremal representations in the Neveu-Schwarz sector.
@@ 1. Introduction
In our paper [8], we gave a detailed proof of the classification of non-extremal (=massive) unitary modules over all unitary minimal simple $W$-algebras, found the necessary conditions of unitarity of the extremal (=massless) ones, and conjectured that all of them are unitary. We also proved that the extremal modules over $N=3$ and $4$ superconformal algebras are indeed unitary, confirming all the results of [2] and [13]. For the bigN=4big\ N=4 our results in [8] and [10] confirmed the conjectures and the results of [4]. In [9] and [10] we have dealt with the Ramond sector, proving necessary conditions for unitarity and a classification of unitary highest weight modules up to the (still conjectural) exactness property of the twisted quantum Hamiltonian reduction functor ([9, Conjecture 9.11]).
In the present paper we provide a proof of [9, Theorem 9.17] in the cases $\mathfrak{g}=psl(2|2)$, $spo(2|2m)$, $D(2,1;a)$, $F(4)$ that does not rely on [9, Conjecture 9.11] (see Theorem 4.5). This result completes the classification of non-extremal unitary representations for all minimal unitary $W$-algebras since the classification for the remaining cases of $\mathfrak{g}=spo(2|2m+1)$ and $G(3)$ is already provided by [9, Corollary 9.5].
@@
The proofs are based on a functor described in Section 2 between the categories of ordinary modules over a vertex algebra and the Ramond twisted modules. Such functor cannot exist for minimal $W$-algebras in the cases $\mathfrak{g}=spo(2|2m+1)$ and $G(3)$ since in these cases the quantum Hamiltonian reduction of a simple module is not simple ([9, Conjecture 9.11]), while it is simple in the Neveu-Schwarz sector, by Arakawa’s theorem.
We believe that our functor is simply the spectral flow, but we do not need this fact and its proof would take us far from the goal of this paper. Evidence that our spectral flow coincides with that defined in [12] is given by the example of the free boson, developed in [8, Section 6].
@@ 4. Spectral flow and unitary highest weight modules
Similarly, a $\sigma_{R}$-twisted highest weight module $L^{R}(\nu,\ell)$ can be unitary only if $\nu\in P^{k}_{+}(R)$ and $\ell\geq A^{R}(k,\nu)$ with $A^{R}(k,\nu)$ given by [9, (6.31)]. Explicit expressions for both $A^{NS}(k,\nu)$ and $A^{R}(k,\nu)$ are given case by case in [8, Section 12] and [9, Section 10] respectively.
@@
We will prove (4.5) by a case by case inspection using the explicit expressions for $A^{NS}(k,\nu)$ and $A^{R}(k,\nu)$ given in [8, Section 12] and [9, Section 10] respectively.
@@ 4.3. $D(2,1,\tfrac{m}{n})\ \text{\rm with $m,n$ coprime}$
Let $L^{R}(\nu,\ell)$ be the irreducible $\sigma^{R}$-twisted highest weight module of highest weight $(\nu,\ell)$. Let $v_{\nu,\ell}\in L^{R}(\nu,\ell)$ be a highest weight vector. Recall that this means that $v_{\nu,\ell}$ is a cyclic vector in $L^{R}(\nu,\ell)$ that satisfies (4.1), (4.2), (4.3). In the following result we use spectral flow to provide a proof of [9, Theorem 9.17] that does not rely on Conjecture 9.11 of [9].
Recall that a weight $\nu\in P^{+}_{k}$ is said to be Ramond extremal (w.r.t. $\rho_{R}$) if $\nu-\rho_{R}\notin P^{+}_{k}$ or $\nu-\rho_{R}$ is extremal (see [9, (9.3)] and [8, Definition 8.7]).
@@
Finally, if $\ell=A^{R}(k,\mu)$ we use the limiting argument given e.g. in [8, Theorem 11.1], [9, Theorem 9.17] to conclude that $L^{R}(\mu,A(k,\mu))$ is unitary as well. ∎
@@ Acknowledgments
The authors would like to thank Dražen Adamović for useful correspondence.
Victor Kac is partially supported by Simons Travel Grant. Pierluigi Möseneder Frajria and Paolo Papi are partially supported by the PRIN project 2022S8SSW2 - Algebraic and geometric aspects of Lie theory - CUP B53D2300942 0006, a project cofinanced by European Union - Next Generation EU fund.

### v2

# Spectral flow and application to unitarity of representations of minimal $W$-algebras

@@ Abstract.
Using spectral flow, we provide a proof of [10, Theorem 9.17] on unitarity of Ramond twisted non-extremal representations of minimal $W$-algebras that does not rely on the still conjectural exactness of the twisted quantum reduction functor (see Conjecture 9.11 of [10]). When $\mathfrak{g}=spo(2|2n)$, $F(4)$, $D(2,1;\frac{m}{n})$, it is also proven that the unitarity of extremal (=massless) representations of the minimal $W$-algebra $W^{k}_{\min}(\mathfrak{g})$ in the Ramond sector is equivalent to the unitarity of extremal representations in the Neveu-Schwarz sector.
@@ 1. Introduction
In our paper [8], we gave a detailed proof of the classification of non-extremal (=massive) unitary modules over all unitary minimal simple $W$-algebras, found the necessary conditions of unitarity of the extremal (=massless) ones, and conjectured that all of them are unitary. We also proved that the extremal modules over $N=3$ and $4$ superconformal algebras are indeed unitary, confirming all the results of [2] and [14]. For the bigN=4big\ N=4 our results in [8] and [9] confirmed the conjectures and the results of [4]. In [10] and [9] we have dealt with the Ramond sector, proving necessary conditions for unitarity and a classification of unitary highest weight modules up to the (still conjectural) exactness property of the twisted quantum Hamiltonian reduction functor ([10, Conjecture 9.11]).
In the present paper we provide a proof of [10, Theorem 9.17] in the cases $\mathfrak{g}=psl(2|2)$, $spo(2|2m)$, $D(2,1;a)$, $F(4)$ that does not rely on [10, Conjecture 9.11] (see Theorem 4.5). This result completes the classification of non-extremal unitary representations for all minimal unitary $W$-algebras since the classification for the remaining cases of $\mathfrak{g}=spo(2|2m+1)$ and $G(3)$ is already provided by [10, Corollary 9.5].
@@
The proofs are based on a functor described in Section 2 between the categories of ordinary modules over a vertex algebra and the Ramond twisted modules. Such functor cannot exist for minimal $W$-algebras in the cases $\mathfrak{g}=spo(2|2m+1)$ and $G(3)$ since in these cases the quantum Hamiltonian reduction of a simple module is not simple ([10, Conjecture 9.11]), while it is simple in the Neveu-Schwarz sector, by Arakawa’s theorem. In the paper [13] it is proved that the above functor is the spectral flow [12]. We thank Haisheng Li for pointing out to us this result.
@@ 2. A functor between categories of positive energy twisted modules
In the next section we will build up a functor between positive energy $W^{k}_{\min}(\mathfrak{g})$-modules mapping untwisted highest weight modules to Ramond twisted highest weight modules. This functor is obtained by applying twice the construction of the contragredient module. It is shown in [13, Section 2] that this functor coincides with the spectral flow defined in [12]. To keep the paper as self-contained as possible and to take care of the modifications needed to work with conjugate dual modules rather than contragredient ones, we provide in this section some details of this construction.
@@ 4. Spectral flow and unitary highest weight modules
Similarly, a $\sigma_{R}$-twisted highest weight module $L^{R}(\nu,\ell)$ can be unitary only if $\nu\in P^{k}_{+}(R)$ and $\ell\geq A^{R}(k,\nu)$ with $A^{R}(k,\nu)$ given by [10, (6.31)]. Explicit expressions for both $A^{NS}(k,\nu)$ and $A^{R}(k,\nu)$ are given case by case in [8, Section 12] and [10, Section 10] respectively.
@@
We will prove (4.5) by a case by case inspection using the explicit expressions for $A^{NS}(k,\nu)$ and $A^{R}(k,\nu)$ given in [8, Section 12] and [10, Section 10] respectively.
@@ 4.3. $D(2,1,\tfrac{m}{n})\ \text{\rm with $m,n$ coprime}$
Let $L^{R}(\nu,\ell)$ be the irreducible $\sigma^{R}$-twisted highest weight module of highest weight $(\nu,\ell)$. Let $v_{\nu,\ell}\in L^{R}(\nu,\ell)$ be a highest weight vector. Recall that this means that $v_{\nu,\ell}$ is a cyclic vector in $L^{R}(\nu,\ell)$ that satisfies (4.1), (4.2), (4.3). In the following result we use spectral flow to provide a proof of [10, Theorem 9.17] that does not rely on Conjecture 9.11 of [10].
Recall that a weight $\nu\in P^{+}_{k}$ is said to be Ramond extremal (w.r.t. $\rho_{R}$) if $\nu-\rho_{R}\notin P^{+}_{k}$ or $\nu-\rho_{R}$ is extremal (see [10, (9.3)] and [8, Definition 8.7]).
@@
Finally, if $\ell=A^{R}(k,\mu)$ we use the limiting argument given e.g. in [8, Theorem 11.1], [10, Theorem 9.17] to conclude that $L^{R}(\mu,A(k,\mu))$ is unitary as well. ∎
@@ Acknowledgments
The authors would like to thank Dražen Adamović and Haisheng Li for important correspondence. Victor Kac is partially supported by Simons Travel Grant. Pierluigi Möseneder Frajria and Paolo Papi are partially supported by the PRIN project 2022S8SSW2 - Algebraic and geometric aspects of Lie theory - CUP B53D2300942 0006, a project cofinanced by European Union - Next Generation EU fund.